Quick-To-Change Company has decided to computerize its accounting system. The company has two alternatives that both entail the use of a computer.

  • It can lease a computer under a three year contract in which the lease payments will be $4,500 each year, payable at the beginning of each year. The lessor will provide all repairs and maintenance under this arrangement.   

  • It can purchase the computer outright for $11,600 in cash. In this case, they will incur the following repair and maintenance costs in addition to the initial purchase price:

  • Paid at the end of year one$400
    Paid at the end of year two$500

 

The computer is expected to have only a three-year useful life because of obsolescence and technological advancements. It will have no salvage value and will be depreciated using the double-declining-balance method.

Quick-To-Change Company's cost of capital is 12%.  Please ignore any income tax effects in this problem.

 

Required:

A.  Prepare a schedule to show the calculation of the net present value of the out-of-pocket costs for the lease alternative. 

B. Prepare a schedule to show the calculation of the net present value of the out-of-pocket costs for the purchase alternative. 

C.  Indicate which of the two alternatives would be best, given the data provided.
Answer #1

Let us have the discount factors

Discount factor for today = 1

Discount factor for the end of year 1 = 1/(1+0.12) = 1/1.12 = 0.893

Discount factor for the end of year 2 = 1/(1+0.12)^2 = 1/1.2544 = 0.797

A. LEASE ALTERNATIVE

Present value of Lease Payment # 1 ($ 4500 * 1) = $ 4500

Present value of Lease Payment # 2 ($ 4500 * 0.893) = $ 4018.5

Present value of Lease Payment # 3 ($ 4500 * 0.797) = $ 3586.5

Net Present value of Out of pocket lease payments   = $ 12,105

B. PURCHASE ALTERNATIVE

Purchase value of Acquisition Cost = $ 11,600

Present value of year one repairs and maintenance ($ 400 * 0.893) = $ 357.2

Present value of year two repairs and maintenance ($ 500 * 0.797) = $ 398.5

Net Present value of Out of pocket payments for the purcahse option = $ 12,355.7

C. Since, the Net present value of the payments under Option A is lower in comparison with option 2. Therefore, Leasing the asset would be a better option instead of purchase.


answered by: bearincar